OpenStudy (anonymous):

solve 8x - (2x + 6)= 18

OpenStudy (anonymous):

\[8x-2x-6=18\] \[6x=24\] \[x=24/6\] \[x=4\]

OpenStudy (aroub):

8x-2x-6=18 6x=18+6 6x=24 x=4

OpenStudy (anonymous):

x = 4x = 4

OpenStudy (anonymous):

Ok, i have no clue what happens this past 2 weeks but lets just all start over again :D